OpenStudy (anonymous):

what is the slope of 3x+5y=5 and -20+12y=0 are they perpendicular, parallel, or neither?

OpenStudy (mathstudent55):

Solve each equation for y. Then you get an equation in the form y = mx + b m, which multiplies x, is the slope. If the slopes are the same and b is different, the lines are parallel. If the slopes multiply to give you -1, the lines are perpendicular. If neither of the above, the lines are neithe perpendicular, nor parallel.

OpenStudy (mathstudent55):

Start with the first equation: 3x + 5y = 5 5y = -3x + 5 y = (-3/5)x + 1 Slope is -3/5, b = 1 (where line crosses y-axis) Now solve the second eqaution for y and post what you get.

OpenStudy (anonymous):

they are perpendicular

OpenStudy (anonymous):

-20x+12y=0 12y=20x+0 12/20=5/3

OpenStudy (mathstudent55):

12y = 20x divide both sides by 12, y = (20/12)x y = (5/3)x Since (5/3)(-3/5) = -1, lines are perpendicular